Solve this system by substitution:
-3x+3y=4 and -x+y=3

Answer by mananth(16946)

-1 x + 1 y = 3 .............1
1 y = 3 + 1 x
/ 1 =
y = 3 + 1 x
3 x + 3 y = 4 .....................2
Substitute y in (2)
3 x + 3 ( 3 + 1 x ) = 4
3 x + 9 + 3 x = 4
3 x 3 x = -5
6 x = -5
/ 6
x= - 5/ 6
Plug the value of x in (1)
-1 x + 1 y = 3
-1 * - 5/ 6 + 1 y = 3
5/ 6 + 1 y = 3
y = 2 1/ 6
y= 13/ 6
m.ananth@hotmail.ca